# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ VII-A Consensus under Ideal Conditions
Figure 5: Depiction of $|\xi|_{\mathcal{A}}$ as a function of continuous time, where $\xi=\phi(t,j)$—the particular solution to the hybrid system $\mathcal{H}$ generated in the nominal simulation. The vertical axis uses a logarithmic scale; the horizontal axis uses a linear scale. Based on the plot, the distance of $\xi$ to $\mathcal{A}$ is exponentially converging to 00.
Figure 6: (Top Left) Depiction of $\|\tilde{\eta}_{p}(\phi(t,j))\|$, for each agent $p\in\mathcal{V}$, versus continuous time. Note that the vertical axis uses a logarithmic scale, while the horizontal axis uses a linear scale. Although noisy, the magnitude of $\tilde{\eta}_{p}$, along the solution $\phi(t,j)$, tends to 00 as $t+j$ approaches infinity. (Bottom Left) Depiction of the trajectories of $\tau_{p}(\phi(t,j))$ for each $p\in\mathcal{V}$. Since $\tau_{p}$ determines when agent $p$ updates the $\eta_{p}$ component of its controller (i.e., when $\tau_{p}=0$) and the timers in $\{\tau_{q}\}_{q\in\mathcal{V}}$ reach 00 at different times, the variables in $\{\eta_{q}\}_{q\in\mathcal{V}}$ are updated intermittently and asynchronously with respect to each other. (Top Right) Depiction of $\|\tilde{\zeta}_{r}(\phi(t,j))\|$, for each $r\in[M^{\ast}]$, versus continuous time. Note that the vertical axis uses a logarithmic scale, while the horizontal axis uses a linear scale. Although noisy, the magnitude of $\tilde{\zeta}_{r}$, along the solution $\phi(t,j)$, tends to 00 as $t+j$ approaches infinity. (Bottom Right) Depiction of the trajectories of $\rho_{r}(\phi(t,j))$ for each $r\in[M^{\ast}]$. Since $\rho_{r}$ determines when the agents in inter-cluster $r$ update the $\zeta_{qr}$ component of their controller, for $q\in\mathcal{V}^{r}$, and the timers in $\{\rho_{s}\}_{s\in\mathcal{V}}$ reach 00 at different times, the variables in $\{\zeta_{qr}\}_{q\in\mathcal{V}^{r}}$ are updated intermittently and asynchronously with respect to the variables in $\{\zeta_{qs}\}_{q\in\mathcal{V}^{s}}$ for $r\neq s\in[M^{\ast}]$. Note that, for each $p\in\mathcal{V}$, the color of $\|\tilde{\eta}_{p}\|$ matches that of $\tau_{p}$. Likewise, for each $r\in[M^{\ast}]$, the color of $\|\tilde{\zeta}_{r}\|$ matches that of $\rho_{r}$.
VII-B Consensus under Measurement Noise
@@
$\delta_{p}$ $=0.07\sin(0.015t)\sin(0.2t)\cdot 1_{m}$ (74)
$\delta_{pr}$ $=0.04\sin(0.095t)\cos(0.4t)\cdot 1_{m},$
@@
To further examine the robustness of the consensus result to measurement noise, $30$ additional simulations using randomly selected clustered networks and perturbation magnitudes were conducted. Nearly all simulation parameters were unchanged, except $K_{u}=1.5\cdot I_{2}$. Figure 9 shows the results of these simulations. In all cases, $\omega$-approximate rendezvous was achieved.
@@
Figure 9: Illustration of $\|x^{\circ}(\phi(t,j))\|$ versus continuous time for $30$ different simulations. Each color corresponds to a unique scenario and a distinct solution $\phi$ to $\mathcal{K}$. Nearly all simulation parameters are unchanged between runs except for the clustered network and perturbation magnitude. In all cases, the MAS drove $\|x^{\circ}(\phi(t,j))\|$ to a small neighborhood of 00.
VIII Conclusion
This paper develops a distributed consensus controller for clustered network of LTI systems using intermittent and asynchronous output feedback. The controller consists of multiple components that are intermittently updated with output information from neighbors of the same cluster and potentially multiple inter-clusters at different rates. Despite the clustered graph structure, which may result in disconnected distanced-based communication graphs as modeled through the timer dynamics, the MAS can still achieve system-wide consensus.
Future work may consider nonlinear, uncertain, and/or heterogeneous agent models. One may examine consensus under path constraints to guarantee safety or other desired behaviors. More advanced observers can also be developed to promote even less frequent communication and/or measurements.
@@ Appendix A Proof of Lemma 1
Since the graph $\mathcal{G}$ is undirected and defined on $N$ nodes, the associated Laplacian matrix $\mathsf{L}\in\mathbb{R}^{N\times N}$ is symmetric, which implies $\mathsf{L}$ is diagonalizable and has real eigenvalues. Further, since $\mathcal{G}$ is static, the diagonalization of $\mathsf{L}$ is fixed. Let $\gamma\triangleq\{w_{1},w_{2},...,w_{N}\}\subset\mathbb{R}^{N}$ and $\{\lambda_{1},\lambda_{2},...,\lambda_{N}\}\subset\mathbb{R}$ denote the eigenvector and eigenvalue sets of $\mathsf{L}$, respectively, where $\lambda_{p}$ is the eigenvalue corresponding to eigenvector $w_{p}$ for each $p\in[N]$. Since $\mathcal{G}$ is connected, $w_{1}\in\text{Span}(1_{N})$, $\lambda_{1}=0$, and, without loss of generality, $\lambda_{1}<\lambda_{2}\leq...\leq\lambda_{N}$. Observe that $\gamma$ is an ordered basis for $\text{Range}(\mathsf{L})$. We can then employ the Gram-Schmidt process on $\gamma$ to construct an orthogonal basis $\beta^{\prime}=\{v_{1}^{\prime},v_{2}^{\prime},...,v_{N}^{\prime}\}$ for $\text{Range}(\mathsf{L})$, where the normalization of the elements in $\beta^{\prime}$ yields an orthonormal basis $\beta$. That is, $\beta=\{v_{1},v_{2},...,v_{N}\}\subset\mathbb{R}^{N}$, where $v_{p}\triangleq v_{p}^{\prime}/\|v_{p}^{\prime}\|$ and $v_{p}\perp v_{q}$ for all distinct $p,q\in[N]$. Observe $\beta^{\prime}$ and $\beta$ are not necessarily unique. Since $w_{1}\in\text{Span}(1_{N})$, $w_{1}=c\cdot 1_{N}$ for some constant $c\in\mathbb{R}\setminus\{0\}$. Since the Gram-Schmidt process allows for $v_{1}^{\prime}\triangleq|w_{1}|$, it follows that $v_{1}=|c|\cdot 1_{N}/\|c\cdot 1_{N}\|=(\sqrt{N}/N)\cdot 1_{N}$.
@@
Using the elements in $\beta$, let $\mathsf{V}^{\prime}\triangleq[v_{1},v_{2},...,v_{N}]\in\mathbb{R}^{N\times N}$. Similarly, let $\mathsf{D}^{\prime}\triangleq\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})\in\mathbb{R}^{N\times N}$. Recall $\lambda_{1}=0$ and $\lambda_{p}>0$ for every $p=2,3,...,N$. Moreover, $(\mathsf{V}^{\prime})^{-1}=(\mathsf{V}^{\prime})^{\top}$ since the columns of $\mathsf{V}^{\prime}$ form an orthonormal set. Using the diagonalization of $\mathsf{L}$, one has that [^9] The Gram-Schmidt process and normalization do not affect the bases of the eigenspaces of $\mathsf{L}$. Hence, $\{w_{p},\lambda_{p}\}_{p\in\mathcal{V}}$ and $\{v_{p},\lambda_{p}\}_{p\in\mathcal{V}}$ yield different diagonalizations of $\mathsf{L}$.
$\mathsf{L}$ $=\mathsf{V}^{\prime}\mathsf{D}^{\prime}(\mathsf{V}^{\prime})^{-1}=\mathsf{V}^{\prime}\mathsf{D}^{\prime}(\mathsf{V}^{\prime})^{\top}$
$=\sum_{p\in[N]}\lambda_{p}v_{p}v_{p}^{\top}=\lambda_{1}v_{1}v_{1}^{\top}+\sum_{p\in[N]\setminus\{1\}}\lambda_{p}v_{p}v_{p}^{\top}$
$=\mathsf{V}\mathsf{D}\mathsf{V}^{\top},$
@@
such that $\mathsf{D}=\text{diag}(\lambda_{2},\lambda_{3},...,\lambda_{N})\in\mathbb{R}^{N-1\times N-1}$ is a positive definite matrix. Recall $\tilde{\mathtt{e}}_{k}$ denotes the $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N-1}$. Since the columns of $\mathsf{V}$ are pairwise orthogonal, one has that $\mathsf{V}^{\top}v_{p}=\tilde{\mathtt{e}}_{p-1}$ for all $p=2,3,...,N$. Therefore,
@@
$\mathsf{V}\mathsf{V}^{\top}$ $=\sum_{p\in[N]\setminus\{1\}}v_{p}v_{p}^{\top}=\sum_{p\in[N]\setminus\{1\}}v_{p}v_{p}^{\top}+v_{1}v_{1}^{\top}-v_{1}v_{1}^{\top}$
$=\sum_{p\in[N]}v_{p}v_{p}^{\top}-\frac{\sqrt{N}}{N}1_{N}\frac{\sqrt{N}}{N}1_{N}^{\top}$
@@
Hence, the desired result follows. $\blacksquare$
@@ Appendix B Proof of Lemma 3
Item 1: The proof is based on [32, Proposition 2.10]. Observe that the flow map $f$ in (24) is globally Lipschitz. Consider the differential equation $\dot{\xi}=f(\xi)$ with initial condition $\phi(0,0)\in C\setminus D$. Then, there exists a nontrivial, maximal solution $\phi$ for $\mathcal{H}$ satisfying the initial condition. Further, under the construction of $\mathcal{H}$, $G(D)\subset C\cup D$ implying that Item (c) in [32, Proposition 2.10] does not occur. Because the single-valued flow map $f$ is Lipschitz continuous, Item (b) in [32, Proposition 2.10] does not occur. Hence, $\phi$ is complete.
Item 2: The following is guided by the proof of Lemma 3.5 in [34]. Fix a maximal solution $\phi$ of $\mathcal{H}$, and let $(t^{\prime},j^{\prime})\in\text{dom }\phi$.
Case I: Suppose $(t,j)\in\text{dom }\phi$ such that $0\leq t-t^{\prime}\leq T_{\min}$ and $0\leq j-j^{\prime}$. With respect to the dynamics in (6) and (7), at most $N+M^{\ast}$ timers may jump twice during a flow interval of length $T_{\min}$. This occurs when $\tau_{p}(t^{\prime},j^{\prime})=0$ with $\tau_{p}^{+}=T_{\min}$ for each $p\in\mathcal{V}$ and $\rho_{r}(t^{\prime},j^{\prime})=0$ with $\rho_{r}^{+}=T_{\min}$ for each $r\in[M^{\ast}]$.101010 Such a reset is possible when $T_{1}^{p}=T_{3}^{r}$ and $T_{2}^{p}=T_{4}^{r}$ for all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$. If a timer has a value in $(0,T_{\min}]$ at time $(t^{\prime},j^{\prime})$, then that timer jumps once during a flow interval of length $T_{\min}$. Hence, the number of jumps between $(t^{\prime},j^{\prime})$ and $(t,j)$ is bounded as $j-j^{\prime}\leq(N+M^{\ast})(\tfrac{t-t^{\prime}}{T_{\min}}+1)$, which implies that
@@
Case II: Suppose $(t,j)\in\text{dom }\phi$ such that $t-t^{\prime}=T_{\max}$ and $0\leq j-j^{\prime}$. With respect to the dynamics in (6) and (7), at least $N+M^{\ast}$ timers must jump once during a flow interval of length $T_{\max}$. This occurs when $\tau_{p}(t^{\prime},j^{\prime})=T_{\max}$ for each $p\in\mathcal{V}$ and $\rho_{r}(t^{\prime},j^{\prime})=T_{\max}$ for each $r\in[M^{\ast}]$.10 If a timer has a value of 00 at time $(t^{\prime},j^{\prime})$ and is reset to some value in $[T_{\min},T_{\max}]$, then that timer may jump at least twice during a flow interval of length $T_{\max}$. Hence, the number of jumps between $(t^{\prime},j^{\prime})$ and $(t,j)$ is bounded as $(N+M^{\ast})\tfrac{t-t^{\prime}}{T_{\max}}\leq j-j^{\prime}$, which implies that
@@
Combining (75) and (76) while setting $(t^{\prime},j^{\prime})=(0,0)$ yields the result in (29). $\blacksquare$
[Uncaptioned image] Federico M. Zegers (Member, IEEE) received a B.S. in Mechanical Engineering in 2016, a B.S. in Mathematics in 2016, an M.S. in Mechanical Engineering in 2019, and a Ph.D. in Mechanical Engineering in 2021 from the University of Florida, Gainesville, FL, USA. He is currently a Senior Professional Staff I within the Air and Missile Defense Sector of Johns Hopkins University Applied Physics Laboratory, Laurel, MD, USA. His research focuses on Lyapunov-based nonlinear and adaptive control, switched and hybrid systems, multi-agent systems, and robotics.
[Uncaptioned image] Sean Phillips (M’11) received his B.S. and M.S. degrees in Mechanical Engineering from the University of Arizona, Tucson, Arizona, in 2011 and 2013, respectively. In 2018, he received a Ph.D. in Computer Engineering from the University of California – Santa Cruz, California. He is a Research Mechanical Engineer and Technology Advisor of the Space Control Branch at the Air Force Research Laboratory in the Space Vehicles Directorate. He also holds the title of Research Assistant Professor (LAT) at the University of New Mexico in Albuquerque, NM. In 2009, he joined the Hybrid Dynamics and Controls Lab where he received a NASA Space Grant in 2010. In 2010, he received an Undergraduate Research Grant from the University of Arizona Honors College. He received a Space Scholars Internship at the Air Force Research Laboratory in Albuquerque N.M. during the summers of 2011, 2012, and 2017. In 2014, he joined the Hybrid Systems Laboratory at the University of California, Santa Cruz. In 2017, he received the Jack Baskin and Peggy Downes-Baskin Fellowship for his research on autonomous networked systems from the Baskin School of Engineering at the University of California Santa Cruz.

### v3

@@ 7.1 Consensus under Ideal Conditions
Figure 5: Depiction of $|\xi|_{\mathcal{A}}$ as a function of continuous time, where $\xi=\phi(t,j)$—the particular solution to the hybrid system $\mathcal{H}$ generated in the nominal simulation. The vertical axis uses a logarithmic scale; the horizontal axis uses a linear scale. Based on the plot, the distance of $\xi$ to $\mathcal{A}$ is exponentially converging to 0.
Figure 6: (Top Left) Depiction of $\|\tilde{\eta}_{p}(\phi(t,j))\|$, for each agent $p\in\mathcal{V}$, versus continuous time. Note that the vertical axis uses a logarithmic scale, while the horizontal axis uses a linear scale. Although noisy, the magnitude of $\tilde{\eta}_{p}$, along the solution $\phi(t,j)$, tends to 0 as $t+j$ approaches infinity. (Bottom Left) Depiction of the trajectories of $\tau_{p}(\phi(t,j))$ for each $p\in\mathcal{V}$. Since $\tau_{p}$ determines when agent $p$ updates the $\eta_{p}$ component of its controller (i.e., when $\tau_{p}=0$) and the timers in $\{\tau_{q}\}_{q\in\mathcal{V}}$ reach 0 at different times, the variables in $\{\eta_{q}\}_{q\in\mathcal{V}}$ are updated intermittently and asynchronously with respect to each other. (Top Right) Depiction of $\|\tilde{\zeta}_{r}(\phi(t,j))\|$, for each $r\in[M^{\ast}]$, versus continuous time. Note that the vertical axis uses a logarithmic scale, while the horizontal axis uses a linear scale. Although noisy, the magnitude of $\tilde{\zeta}_{r}$, along the solution $\phi(t,j)$, tends to 0 as $t+j$ approaches infinity. (Bottom Right) Depiction of the trajectories of $\rho_{r}(\phi(t,j))$ for each $r\in[M^{\ast}]$. Since $\rho_{r}$ determines when the agents in inter-cluster $r$ update the $\zeta_{qr}$ component of their controller, for $q\in\mathcal{V}^{r}$, and the timers in $\{\rho_{s}\}_{s\in\mathcal{V}}$ reach 0 at different times, the variables in $\{\zeta_{qr}\}_{q\in\mathcal{V}^{r}}$ are updated intermittently and asynchronously with respect to the variables in $\{\zeta_{qs}\}_{q\in\mathcal{V}^{s}}$ for $r\neq s\in[M^{\ast}]$. Note that, for each $p\in\mathcal{V}$, the color of $\|\tilde{\eta}_{p}\|$ matches that of $\tau_{p}$. Likewise, for each $r\in[M^{\ast}]$, the color of $\|\tilde{\zeta}_{r}\|$ matches that of $\rho_{r}$.
7.2 Consensus under Measurement Noise
@@
$\delta_{p}=0.07\sin(0.015t)\sin(0.2t)\cdot 1_{m}\quad\delta_{pr}=0.04\sin(0.095t)\cos(0.4t)\cdot 1_{m},$ (75)
@@
To further examine the robustness of the consensus result to measurement noise, $30$ additional simulations using randomly selected clustered networks and perturbation magnitudes were conducted. Nearly all simulation parameters were unchanged, except $K_{u}=1.5\cdot I_{2}$. Figure 9 conveys the results of these simulations. In all cases, $\omega$-approximate rendezvous was achieved.
@@
Figure 9: Illustration of $\|x^{\circ}(\phi(t,j))\|$ versus continuous time for $30$ different simulations. Each color corresponds to a unique scenario and a distinct solution $\phi$ to $\mathcal{K}$. Nearly all simulation parameters are unchanged between runs except for the clustered network and perturbation magnitude. In all cases, the MAS drove $\|x^{\circ}(\phi(t,j))\|$ to a small neighborhood of 0.
8 Conclusion
This paper develops a distributed consensus controller that is hybrid in structure for a clustered network of identical LTI systems utilizing intermittent and asynchronous output feedback. Specifically, the controller of each agent consists of multiple components that are intermittently updated at different rates with output information from same-cluster neighbors and same-inter-cluster neighbors. Although the communication network is modeled as a static, connected, and undirected graph, the flow out output information over the edges of this graph can be intermittent and asynchronous.
Future work may consider nonlinear, uncertain, and/or heterogeneous agent models as well as directed communication graphs given their practical importance. One way of extending the result to heterogeneous agent dynamics is by approximating each dynamics with a common LTI system and then designing the controller of each agent to either adaptively or robustly compensate for the disturbance defined by the difference between the LTI system and the true agent dynamics. One may examine consensus under path constraints to guarantee safety or other desired behaviors. More advanced estimators can also be developed to promote less frequent communication and/or measurements. Of particular importance is the derivation of a constructive extension of Corollary 1 to simplify the design of the proposed consensus controller. Moreover, the effect of the clustered network structure on the solution space of $\mathsf{M}(\tau,\rho)<\mathbf{0}$ should be investigated.
@@ Appendix A Proof of Lemma 1
Since the graph $\mathcal{G}$ is undirected and defined on $N$ nodes, the associated Laplacian matrix $\mathsf{L}\in\mathbb{R}^{N\times N}$ is symmetric, which implies $\mathsf{L}$ is diagonalizable and has real eigenvalues. Further, since $\mathcal{G}$ is static, the diagonalization of $\mathsf{L}$ is fixed. Let $\gamma\coloneqq\{w_{1},w_{2},...,w_{N}\}\subset\mathbb{R}^{N}$ and $\{\lambda_{1},\lambda_{2},...,\lambda_{N}\}\subset\mathbb{R}$ denote the eigenvector and eigenvalue sets of $\mathsf{L}$, respectively, where $\lambda_{p}$ is the eigenvalue corresponding to eigenvector $w_{p}$ for each $p\in[N]$. Since $\mathcal{G}$ is connected, $w_{1}\in\text{Span}(1_{N})$, $\lambda_{1}=0$, and, without loss of generality, $\lambda_{1}<\lambda_{2}\leq...\leq\lambda_{N}$. Observe that $\gamma$ is an ordered basis for $\text{Range}(\mathsf{L})$. We can then employ the Gram-Schmidt process on $\gamma$ to construct an orthogonal basis $\beta^{\prime}=\{v_{1}^{\prime},v_{2}^{\prime},...,v_{N}^{\prime}\}$ for $\text{Range}(\mathsf{L})$, where the normalization of the elements in $\beta^{\prime}$ yields an orthonormal basis $\beta$. That is, $\beta=\{v_{1},v_{2},...,v_{N}\}\subset\mathbb{R}^{N}$, where $v_{p}\coloneqq v_{p}^{\prime}/\|v_{p}^{\prime}\|$ and $v_{p}\perp v_{q}$ for all distinct $p,q\in[N]$. Observe $\beta^{\prime}$ and $\beta$ are not necessarily unique. Since $w_{1}\in\text{Span}(1_{N})$, $w_{1}=c\cdot 1_{N}$ for some constant $c\in\mathbb{R}\setminus\{0\}$. Since the Gram-Schmidt process allows for $v_{1}^{\prime}\coloneqq|w_{1}|$, it follows that $v_{1}=|c|\cdot 1_{N}/\|c\cdot 1_{N}\|=(\sqrt{N}/N)\cdot 1_{N}$.
@@
Using the elements in $\beta$, let $\mathsf{V}^{\prime}\coloneqq[v_{1},v_{2},...,v_{N}]\in\mathbb{R}^{N\times N}$. Similarly, let $\mathsf{D}^{\prime}\coloneqq\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})\in\mathbb{R}^{N\times N}$. Recall $\lambda_{1}=0$ and $\lambda_{p}>0$ for every p=2,3,,Np=2,3,...,N. Moreover, $(\mathsf{V}^{\prime})^{-1}=(\mathsf{V}^{\prime})^{\top}$ since the columns of $\mathsf{V}^{\prime}$ form an orthonormal set. Using the diagonalization of $\mathsf{L}$, one has that [^12] The Gram-Schmidt process and normalization do not affect the bases of the eigenspaces of $\mathsf{L}$. Hence, $\{w_{p},\lambda_{p}\}_{p\in\mathcal{V}}$ and $\{v_{p},\lambda_{p}\}_{p\in\mathcal{V}}$ yield different diagonalizations of $\mathsf{L}$.
$\mathsf{L}=\mathsf{V}^{\prime}\mathsf{D}^{\prime}(\mathsf{V}^{\prime})^{-1}=\mathsf{V}^{\prime}\mathsf{D}^{\prime}(\mathsf{V}^{\prime})^{\top}=\sum_{p\in[N]}\lambda_{p}v_{p}v_{p}^{\top}=\lambda_{1}v_{1}v_{1}^{\top}+\sum_{p\in[N]\setminus\{1\}}\lambda_{p}v_{p}v_{p}^{\top}=\mathsf{V}\mathsf{D}\mathsf{V}^{\top},$
@@
such that $\mathsf{D}=\text{diag}(\lambda_{2},\lambda_{3},...,\lambda_{N})\in\mathbb{R}^{N-1\times N-1}$ is a positive definite matrix. Recall $\tilde{\mathtt{e}}_{k}$ denotes the $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N-1}$. Since the columns of $\mathsf{V}$ are pairwise orthogonal, one has that $\mathsf{V}^{\top}v_{p}=\tilde{\mathtt{e}}_{p-1}$ for all p=2,3,,Np=2,3,...,N. Therefore,
@@
$\mathsf{V}\mathsf{V}^{\top}$ $=\sum_{p\in[N]\setminus\{1\}}v_{p}v_{p}^{\top}=\sum_{p\in[N]\setminus\{1\}}v_{p}v_{p}^{\top}+v_{1}v_{1}^{\top}-v_{1}v_{1}^{\top}=\sum_{p\in[N]}v_{p}v_{p}^{\top}-\frac{\sqrt{N}}{N}1_{N}\frac{\sqrt{N}}{N}1_{N}^{\top}$
@@
Hence, the desired result follows. ∎
@@ Appendix B Proof of Lemma 3
Item 1: The proof is based on [Goebel.Sanfelice.ea2012, Proposition 2.10]. Consider the differential equation $\dot{\xi}=f(\xi)$ with initial condition $\phi(0,0)\in C\setminus D$. Since the flow map $f$ in (25) is globally Lipschitz, there exists a nontrivial, maximal solution $\phi$ for $\mathcal{H}$ satisfying the initial condition. Further, under the construction of $\mathcal{H}$, $G(D)\subset C\cup D$ implying that Item (c) in [Goebel.Sanfelice.ea2012, Proposition 2.10] does not occur. Since the single-valued flow map $f$ is Lipschitz continuous, Item (b) in [Goebel.Sanfelice.ea2012, Proposition 2.10] does not occur. Hence, $\phi$ is complete.
Item 2: The following is guided by the proof of Lemma 3.5 in [Li.Phillips.ea2018]. Fix a maximal solution $\phi$ of $\mathcal{H}$, and let $(t^{\prime},j^{\prime})\in\text{dom }\phi$.
Case I: Suppose $(t,j)\in\text{dom }\phi$ such that $0\leq t-t^{\prime}\leq T_{\min}$ and $0\leq j-j^{\prime}$. With respect to the dynamics in (8) and (9), at most $N+M^{\ast}$ timers may jump twice during a flow interval of length $T_{\min}$. This occurs when $\tau_{p}(t^{\prime},j^{\prime})=0$ with $\tau_{p}^{+}=T_{\min}$ for each $p\in\mathcal{V}$ and $\rho_{r}(t^{\prime},j^{\prime})=0$ with $\rho_{r}^{+}=T_{\min}$ for each $r\in[M^{\ast}]$. [^13] Such a reset is possible when $T_{1}^{p}=T_{3}^{r}$ and $T_{2}^{p}=T_{4}^{r}$ for all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$. If a timer has a value in $(0,T_{\min}]$ at time $(t^{\prime},j^{\prime})$, then that timer jumps once during a flow interval of length $T_{\min}$. Hence, the number of jumps between $(t^{\prime},j^{\prime})$ and $(t,j)$ is bounded as $j-j^{\prime}\leq(N+M^{\ast})(\tfrac{t-t^{\prime}}{T_{\min}}+1)$, which implies that
@@
Case II: Suppose $(t,j)\in\text{dom }\phi$ such that $t-t^{\prime}=T_{\max}$ and $0\leq j-j^{\prime}$. With respect to the dynamics in (8) and (9), at least $N+M^{\ast}$ timers must jump once during a flow interval of length $T_{\max}$. This occurs when $\tau_{p}(t^{\prime},j^{\prime})=T_{\max}$ for each $p\in\mathcal{V}$ and $\rho_{r}(t^{\prime},j^{\prime})=T_{\max}$ for each $r\in[M^{\ast}]$.13 If a timer has a value of 0 at time $(t^{\prime},j^{\prime})$ and is reset to some value in $T_{\min},T_{\max}]$, then that timer may jump at least twice during a flow interval of length $T_{\max}$. Hence, the number of jumps between $(t^{\prime},j^{\prime})$ and $(t,j)$ is bounded as $(N+M^{\ast})\tfrac{t-t^{\prime}}{T_{\max}}\leq j-j^{\prime}$, which implies that
@@
Combining (76) and (77) while setting $(t^{\prime},j^{\prime})=(0,0)$ yields the result in (30). Lastly, maximal solutions of $\mathcal{H}$ being complete and the inequality in (30) imply $\phi$ is non-Zeno.